OpenStudy (anonymous):

Find X and Y y=2x-3 y=x-1

OpenStudy (nurali):

y=2x-3---(1) y=x-1----(2) In equation (1) put 'y' in equation (2) y=x-1 2x-3=x-1 2x-x=-1+3 x=2 put 'x' in equation (1) y=2x-3 y=2(2)-3 y=1 (x,y)=(2,1)

OpenStudy (anonymous):

Thanks:)

OpenStudy (nurali):

My Pleasure.